# RayFormer: Improving Query-Based Multi-Camera 3D Object Detection via Ray-Centric Strategies

### v2

@@ Abstract.
The recent advances in query-based multi-camera 3D object detection are featured by initializing object queries in the 3D space, and then sampling features from perspective-view images to perform multi-round query refinement. In such a framework, query points near the same camera ray are likely to sample similar features from very close pixels, resulting in ambiguous query features and degraded detection accuracy. To this end, we introduce RayFormer, a camera-ray-inspired query-based 3D object detector that aligns the initialization and feature extraction of object queries with the optical characteristics of cameras. Specifically, RayFormer transforms perspective-view image features into bird’s eye view (BEV) via the lift-splat-shoot method and segments the BEV map to sectors based on the camera rays. Object queries are uniformly and sparsely initialized along each camera ray, facilitating the projection of different queries onto different areas in the image to extract distinct features. Besides, we leverage the instance information of images to supplement the uniformly initialized object queries by further involving additional queries along the ray from 2D object detection boxes. To extract unique object-level features that cater to distinct queries, we design a ray sampling method that suitably organizes the distribution of feature sampling points on both images and bird’s eye view. Extensive experiments are conducted on the nuScenes dataset to validate our proposed ray-inspired model design. The proposed RayFormer achieves 55.5% mAP and 63.3% NDS, respectively. Our codes will be made available.
@@ 1. Introduction
  • We conduct experiments on the challenging nuScenes dataset. Under the same backbone and training configuration as other studies, RayFormer achieves state-of-the-art performance on the test set, with mAP at 55.5% and NDS at 63.4%.
@@ 2. Related work
BEV-based 3D Object Detection: Numerous BEV-based detection methods are currently available, primarily distinguished by their view transformation method - either forward projection (Zhou et al., 2023; Zhang et al., 2022) or backward projection (Yang et al., 2023; Chu et al., 2023). BEVDet (Huang et al., 2021) represents a pioneering effort in employing the LSS method to project image features forward onto the BEV, which is built using existing modules and enhanced by a unique data augmentation strategy and an improved non-maximum suppression technique. Building upon BEVDet, BEVDepth (Li et al., 2023a) introduces the supervision for depth estimation and designs a depth refinement module. BEVFormer (Li et al., 2022c) divides the BEV map into grids and lifts them to pillars, selecting specific 3D positions for projection and sampling on the image features. It leverages deformable cross-attention (Zhu et al., 2021) to extract spatial features across camera views, constructing BEV features and incorporating temporal data for enhanced detection (Park et al., 2023; Li et al., 2022a; Huang and Huang, 2022). FB-BEV (Li et al., 2023b) introduces a forward-backward view transformation module, enhancing both methods to yield improved BEV representations. HOP (Zong et al., 2023) designs short-term, long-term temporal decoders, and an additional object decoder to predict objects using generated pseudo BEV features, forcing the detector to capture both the spatial and temporal motion of objects.
@@
Query-based 3D Object Detection: DETR3D (Wang et al., 2021) and PETR (Liu et al., 2022b) employ the transformer decoder to interpret the extracted image features. The former projects sparse 3D object queries onto images to index 2D features, while the latter integrates the positional embedding of 3D coordinates into image features, yielding 3D position-aware features. StreamPETR (Wang et al., 2023c) extends the PETR series with a novel object-centric temporal mechanism for long-sequence modeling. It operates online, leveraging long-term historical data through frame-by-frame object query propagation. MV2D (Wang et al., 2023b) enhances multi-camera 3D object detection by leveraging 2D object detectors to generate object-specific queries based on image semantics. Sparse4D (Lin et al., 2022) meticulously refines anchor boxes through sparse amalgamation and sampling of spatial-temporal features. It assigns several 4D keypoints to each 3D anchor and fuses these features hierarchically across multiple views and scales to yield high-caliber instance features. SparseBEV (Liu et al., 2023a) introduces a fully sparse 3D object detection framework that fuses scale-adaptive attention and integrates adaptive spatio-temporal sampling and mixing.
@@
Polar Representation for 3D Perception: The majority of 3D perception studies based on polar representation construct the BEV features using polar coordinates rather than Cartesian coordinates. PolarNet (Zhang et al., 2020) introduces a LiDAR-specific segmentation algorithm that uses a polar BEV representation, rather than traditional spherical or square BEV, thereby balancing point distribution across polar grid cells. PolarBEV (Liu et al., 2022a) introduces polar embedding decomposition and rearranges polar grids into an array-like structure for efficient processing, updating the BEV surface based on a hypothetical plane and using height-based feature transformation. PolarFormer (Jiang et al., 2022) designs a cross-attention-based polar detection head and a multi-scale polar representation learning strategy to handle irregular polar BEV grids and varying object scales. TaDe (Zhao et al., 2024) presents a two-step approach that uses an autoencoder to reconstruct noisy BEV maps, optimizes feature correlations, and converts BEV maps into polar coordinates for enhanced alignment. Contrary to these methods, we construct an implicit query-based polar representation, instead of the BEV feature, through sparse radial query initialization, complemented with a ray sampling method that aligns with the radial query distribution.
@@ 3.4. Ray Sampling
where $F^{b}_{t,j}|j\in\{1,2,...,N_{s}\}\}$ denotes the BEV feature point at one of the $N_{s}$ sampling locations of the t-th frame with the weight $w_{t}$.
@@
In addition to the uniformly distributed base queries described in Sec 3.3, we use 2D object detection to guide the initialization of foreground queries, which are then merged with the base queries for prediction. As shown in Fig. 4(a), using the camera intrinsics, the 2D points projected from the 3D positions on the same BEV ray fall on an almost vertical line, allowing us to judge the presence of a 3D object on a given BEV ray by ascertaining if any point on the ray can be projected onto an object’s vertical region in the image. To do this, we extend the detected 2D bounding boxes to the full height of the image to define a broad foreground region. As a result, on the BEV map, we pre-define uniformly distributed rays of varying densities according to the prior dimensions of different categories. We select rays whose midpoints can be projected into the foreground region of each category as the foreground rays. Thus, without relying on the accuracy of depth estimation, the results of 2D object detection alone can guide the initialization of additional queries on these foreground rays. This increases the density of queries around real objects.
@@
Sparse4Dv2 (Lin et al., 2023) 704$\times$256 ResNet50 100 0.439 0.539 0.598 0.270 0.475 0.282 0.179
@@ 4.2. Implementation Details
RayFormer is trained using the AdamW (Loshchilov and Hutter, 2019) optimizer with a global batch size of 8. The initial learning rates for the backbone and other parameters are set at 2e-5 and 2e-4, respectively, with a cosine annealing (Loshchilov and Hutter, 2017) decay strategy. We employ standard 2D feature encoding networks, ResNet (He et al., 2016) and VoVNet-99 (V2-99) (Lee et al., 2019), as our backbones. Following previous methodologies (Liu et al., 2023a; Lin et al., 2022; Wang et al., 2023c), the parameters of ResNet are pretrained on the nuImages (Caesar et al., 2020) dataset, while the parameters of V2-99 are pretrained from DD3D (Park et al., 2021) with extra data. All models undergo training for 24 epochs unless stated otherwise. Neither CBGS (Zhu et al., 2019b) nor test time augmentation are utilized in any of the experiments.
@@
BEVFormer(Li et al., 2022c) 704$\times$256 0.297 0.379 0.739
BEVDet(Huang et al., 2021) 704$\times$256 0.298 0.379 0.725
@@
BEVDepth(Li et al., 2023a) 704$\times$256 0.322 0.367 0.707
Sparse4D(Lin et al., 2022) 704$\times$256 0.322 0.401 0.747
@@ 4.3. Main Results
Comparison with the State-of-the-art Methods: Among all the non-temporal models, RayFormer achieves the highest scores in mAP and NDS on the nuScenes validation set, showing significant improvements of 2.1% and 0.4% respectively over SparseBEV, as illustrated in Tab. 3. The comparison results with multi-frame input (T=8) on the val set are presented in Tab. 1. With an input resolution of $704\times 256$ and a ResNet50 backbone, RayFormer reaches 45.9% in mAP and 55.8% in NDS, outperforming StreamPETR by 0.9% and 0.8% respectively. At an input resolution of $512\times 1408$ and a ResNet101 backbone, RayFormer scores 51.1% in mAP and 59.4% in NDS, surpassing StreamPETR by 0.7% and 0.2%, respectively. The comparison of RayFormer with other advanced methods on the nuScenes test set is presented in Tab. 2. For this test, we employ an input resolution of $1600\times 640$ and selected VoVNet-99 as the backbone. RayFormer scores 55.5% in mAP and 63.3% in NDS, which surpasses the baseline SparseBEV by 1.2% and 0.6% respectively.
@@
Visualization Results:
Fig. 5 presents a visualization example from our system, utilizing an input resolution of $512\times 1408$ and ResNet101 as the backbone. As depicted on the right, the BEV layout presents the entire scene, where predictions and ground truths are denoted by blue and green boxes respectively. The left displays projections of detected 3D bounding boxes onto the image plane, color-coding cars, pedestrians, and bicycles in yellow, blue, and red respectively. Notably, missed detections are circled in dashed red, corresponding to their counterparts on the right. The visualization of ray sampling method is shown in Fig. 6, which denotes sampling points of high-scoring queries on both BEV plane and images with red dots. Owing to the ray sampling strategy, the sampling points show a radial pattern on the BEV and a vertical pattern in images, predominantly projecting onto the object areas.
@@ 4.4. Ablation Studies
Initialization and Sampling Configuration: In Tab. 6, we streamline our RayFormer by omitting the 2D auxiliary learning task and 2D guided foreground query supplement, offering a more straightforward comparison. The first four rows display the outcomes varying by the number of rays and depth segments during query initialization. Upon comparing (A)-(C), there is an improvement in detection performance as more rays are set. Comparing (B) and (D), initiating 6 queries on each ray yields mAP and NDS that are 0.9% and 1.0% higher respectively than with 3 queries. Altering the number of adaptive ray points per query, as seen in (B) and (E), does not cause significant differences. By contrast, increasing the number of sampling offsets per ray point from 1 to 4 (B and F), results in a 1.1% boost in both mAP and NDS. For an optimal balance between precision and efficiency, we chose (B) as the default configuration.
Number of Historical Frames: The detection performance is considerably influenced by multi-frame input, as illustrated in Tab. 7. By incorporating just one additional historical frame in addition to the key-frame input, we see a 4.1% rise in mAP and a 7.4% surge in NDS, particularly due to a substantial drop in mAVE. As we further increase the number of historical frame inputs, we observe consistent decreases in mAVE from 0.800 to 0.267, and mATE from 0.709 to 0.601. Thus, it underscores the significant role that sequential frame input plays in velocity prediction, leading to enhanced accuracy in object localization and recognition.

### v4

@@ Abstract.
The recent advances in query-based multi-camera 3D object detection are featured by initializing object queries in the 3D space, and then sampling features from perspective-view images to perform multi-round query refinement. In such a framework, query points near the same camera ray are likely to sample similar features from very close pixels, resulting in ambiguous query features and degraded detection accuracy. To this end, we introduce RayFormer, a camera-ray-inspired query-based 3D object detector that aligns the initialization and feature extraction of object queries with the optical characteristics of cameras. Specifically, RayFormer transforms perspective-view image features into bird’s eye view (BEV) via the lift-splat-shoot method and segments the BEV map to sectors based on the camera rays. Object queries are uniformly and sparsely initialized along each camera ray, facilitating the projection of different queries onto different areas in the image to extract distinct features. Besides, we leverage the instance information of images to supplement the uniformly initialized object queries by further involving additional queries along the ray from 2D object detection boxes. To extract unique object-level features that cater to distinct queries, we design a ray sampling method that suitably organizes the distribution of feature sampling points on both images and bird’s eye view. Extensive experiments are conducted on the nuScenes dataset to validate our proposed ray-inspired model design. The proposed RayFormer achieves superior performance of 55.5% mAP and 63.3% NDS, respectively.
3D Object Detection, Multiple Camera, Autonomous Driving
journalyear: 2024copyright: acmlicensedconference: Proceedings of the 32nd ACM International Conference on Multimedia; October 28-November 1, 2024; Melbourne, VIC, Australiabooktitle: Proceedings of the 32nd ACM International Conference on Multimedia (MM ’24), October 28-November 1, 2024, Melbourne, VIC, Australiadoi: 10.1145/3664647.3681103isbn: 979-8-4007-0686-8/24/10ccs: Computing methodologies Object detection
@@ 1. Introduction
  • We conduct experiments on the challenging nuScenes dataset. Under the same backbone and training configuration as other studies, RayFormer achieves state-of-the-art performance on the test set, with mAP at 55.5% and NDS at 63.3%.
@@ 2. Related work
BEV-based 3D Object Detection: BEV-based 3D Object Detection has seen various innovative approaches, mainly categorized by their view transformation techniques—either through forward projection (Zhou et al., 2023; Zhang et al., 2022), which pioneers the use of the LSS method for feature projection onto BEV, or backward projection techniques (Yang et al., 2023; Chu et al., 2023). BEVDet (Huang et al., 2021) is distinguished by its unique data augmentation and refined non-maximum suppression. BEVDepth (Li et al., 2023a) extends this foundation by incorporating depth estimation supervision and a depth refinement module. BEVFormer (Li et al., 2022c) further advances the field by dividing the BEV into grids, employing deformable cross-attention (Zhu et al., 2021) to construct BEV features with temporal integration for enhanced detection capabilities (Park et al., 2023; Li et al., 2022a; Huang and Huang, 2022). FB-BEV (Li et al., 2023b) enhances these techniques by introducing a forward-backward view transformation module, while HOP (Zong et al., 2023) innovates with temporal decoders and object prediction using pseudo BEV features, capturing both spatial and temporal aspects of object motion.
@@
Query-based 3D Object Detection: DETR3D (Wang et al., 2021) and PETR (Liu et al., 2022b) utilize transformer decoders to process image features, where the former projects 3D object queries onto 2D features, and the latter integrates the positional embedding of 3D coordinates into image features, yielding 3D position-aware features. StreamPETR (Wang et al., 2023c) builds on PETR with an object-centric temporal mechanism for long-sequence modeling, utilizing historical data through object query propagation. MV2D (Wang et al., 2023b) improves multi-camera detection by generating queries from 2D object detections based on image semantics. Sparse4D (Lin et al., 2022) meticulously refines anchor boxes through sparse amalgamation and sampling of spatial-temporal features, assigning multiple 4D keypoints to each 3D anchor and fusing these features hierarchically across different views and scales. SparseBEV (Liu et al., 2023a) introduces a fully sparse 3D object detection framework that fuses scale-adaptive attention and integrates adaptive spatio-temporal sampling and mixing.
@@
Polar Representation for 3D Perception: In the realm of 3D perception, studies leveraging polar representation have shifted from Cartesian to polar coordinates for BEV feature construction. PolarNet (Zhang et al., 2020) pioneers a segmentation approach using a polar BEV representation, optimizing point distribution across grid cells. PolarBEV (Liu et al., 2022a) enhances processing efficiency through polar embedding decomposition and height-based feature transformation on a hypothetical plane. PolarFormer (Jiang et al., 2022) addresses irregular polar BEV grids and varying object scales with a cross-attention-based detection head and multi-scale learning strategy. TaDe (Zhao et al., 2024) presents a two-step approach that uses an autoencoder to reconstruct noisy BEV maps, optimizes feature correlations, and converts BEV maps into polar coordinates for enhanced alignment. Contrary to these methods, we construct an implicit query-based polar representation, instead of the BEV feature, through sparse radial query initialization, complemented with a ray sampling method that aligns with the radial query distribution.
@@ 3.4. Ray Sampling
where $\{F^{b}_{t,j}|j\in\{1,2,...,N_{s}\}\}$ denotes the BEV feature points at $N_{s}$ sampling locations of the t-th frame with the weight $w_{t}$.
@@
In addition to the uniformly distributed base queries described in Sec 3.3, we use 2D object detection to guide the initialization of foreground queries, which are then merged with the base queries for prediction. As shown in Fig. 4(a), by utilizing camera intrinsics, we observe that 2D projections from 3D positions on a BEV ray form a near-vertical line, facilitating the detection of 3D objects by checking if any point on the ray aligns with the object’s vertical image region. To do this, we extend the detected 2D bounding boxes to the full height of the image to define a broad foreground region. As a result, on the BEV map, we pre-define uniformly distributed rays of varying densities according to the prior dimensions of different categories. We select rays whose midpoints can be projected into the foreground region of each category as the foreground rays. Thus, without relying on the accuracy of depth estimation, the results of 2D object detection alone can guide the initialization of additional queries on these foreground rays. This increases the density of queries around real objects.
@@
Sparse4Dv2 (Lin et al., 2023) 256$\times$704 ResNet50 100 0.439 0.539 0.598 0.270 0.475 0.282 0.179
@@ 4.2. Implementation Details
RayFormer is trained using the AdamW (Loshchilov and Hutter, 2019) optimizer with a global batch size of 8. The initial learning rates for the backbone and other parameters are set at 2e-5 and 2e-4, respectively, with a cosine annealing (Loshchilov and Hutter, 2017) decay strategy. We utilize ResNet (He et al., 2016) and VoVNet-99 (V2-99) (Lee et al., 2019) as our backbone networks for feature encoding. ResNet parameters are pretrained on nuImages (Caesar et al., 2020), and V2-99 parameters are initialized from DD3D (Park et al., 2021). All models are trained for 24 epochs without the use of CBGS (Zhu et al., 2019b) or test time augmentation.
@@
BEVFormer(Li et al., 2022c) 256$\times$704 0.297 0.379 0.739
BEVDet(Huang et al., 2021) 256$\times$704 0.298 0.379 0.725
@@
BEVDepth(Li et al., 2023a) 256$\times$704 0.322 0.367 0.707
Sparse4D(Lin et al., 2022) 256$\times$704 0.322 0.401 0.747
@@ 4.3. Main Results
Comparison with the State-of-the-art Methods: Among all the non-temporal models, RayFormer achieves the highest scores in mAP and NDS on the nuScenes validation set, showing significant improvements of 2.1% and 0.4% respectively over SparseBEV, as illustrated in Tab. 3. The comparison results with multi-frame input (T=8) on the val set are presented in Tab. 1. With an input resolution of $256\times 704$ and a ResNet50 backbone, RayFormer reaches 45.9% in mAP and 55.8% in NDS, outperforming StreamPETR by 0.9% and 0.8% respectively. At an input resolution of $512\times 1408$ and a ResNet101 backbone, RayFormer scores 51.1% in mAP and 59.4% in NDS, surpassing StreamPETR by 0.7% and 0.2%, respectively. The comparison of RayFormer with other advanced methods on the nuScenes test set is presented in Tab. 2. For this test, we employ an input resolution of $640\times 1600$ and selected VoVNet-99 as the backbone. RayFormer scores 55.5% in mAP and 63.3% in NDS, which surpasses the baseline SparseBEV by 1.2% and 0.6% respectively.
@@
Visualization Results: Fig. 5 showcases a system visualization with an input resolution of $512\times 1408$, using ResNet101 as the backbone. The BEV layout on the right captures the full scene, with blue and green boxes indicating predictions and ground truths. The left side shows 3D bounding box projections onto the image plane, with yellow, blue, and red coding for cars, pedestrians, and bicycles, respectively. Missed detections are outlined in dashed red, matching the BEV view. Fig. 6 illustrates the ray sampling method, marking high-scoring query sampling points on the BEV plane and images with red dots. These points exhibit a radial pattern on the BEV and a vertical alignment in images, focusing on object areas due to the radial sampling strategy.
@@ 4.4. Ablation Studies
Initialization and Sampling Configuration: In Tab. 6, we simplify our RayFormer by excluding the 2D auxiliary task and 2D-guided query enhancement, focusing on the core performance. The table shows how detection results vary with the number of rays and depth segments in query initialization. Increasing the number of rays improves performance, as seen in the progression from (A) to (C). Specifically, using 6 queries per ray (D) over 3 queries (B) increases mAP and NDS by 0.9% and 1.0%, respectively. Adjusting the adaptive ray points per query, as in (B) versus (E), has minimal impact. However, increasing the sampling offsets per ray point from 1 to 4, as in (B) versus (F), significantly enhances mAP and NDS by 1.1%. We select configuration (B) as the default for its optimal balance of precision and efficiency.
Number of Historical Frames: The impact of multi-frame input on detection performance is evident in Tab. 7. Adding a single historical frame to the key-frame input results in a notable increase in mAP by 4.1% and a significant boost in NDS by 7.4%, primarily due to a marked reduction in mAVE. Continued increase in the number of historical frames further reduces mAVE and mATE, highlighting the crucial contribution of sequential inputs to velocity estimation and the overall improvement in object detection accuracy.
@@ 5. Conclusion
Acknowledgements.
This work is supported by the National Natural Science Foundation of China (No. 62332016).